OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. c times the square root of c + 12 times c times the square root of c − 2 times c times the square root of c 11c c times the square root of c 11 times c times the square root of c 11 the square root of c

OpenStudy (dumbcow):

\[c \sqrt{c} +12c \sqrt{c} -2c \sqrt{c}\] These are like terms, they all have a csqrt(c) just add the coefficients its the same as if it was , x+12x-2x

OpenStudy (anonymous):

so would it be 11 is the square root of c?